OpenStudy (anonymous):

Given: x + y = 6. If x ∈ {-15, -6, -1}, then which of the following sets of ordered pairs are solutions? {(-15, -21), (-6, -12), (-1, -7)} {(-15, 21), (-6, 12), (-1, 7)} {(-15, 21), (-6, 12), (-1, -7)}

OpenStudy (anonymous):

To solve this, you'll want to plug each element of the set into the equation as x and solve each for y. That will give you three pairs of x and y values.

OpenStudy (campbell_st):

rewrite the equation y = 6 - x substitiute x = -15 to find y = 6 - (-15) y = 21 then x = -6 to find another y lastly x = - 1 to find the last y this will give the solution set of ordered pairs

OpenStudy (anonymous):

second one